OpenStudy (anonymous):

given the function f(x)=x^3+x^2-2x+1 what is resulting when f(x) is shifted to the left 1 unit?

OpenStudy (mikeyymonroe):

sorry hun i will try my hardest to help but without my calculator i cant know hat the function looks like so like i said ill try. :)

OpenStudy (anonymous):

its Ok :)

OpenStudy (abdullah_me):

f(x)=(x+1)^3+ (x+1)^2-2(x+1)+1

OpenStudy (anonymous):

Thank you so much do you think you can help me with another one I only have on more question I am stuck on

OpenStudy (abdullah_me):

ok

OpenStudy (anonymous):

Given the function f(x) = −3x3 + 9x2 − 2x + 3, what part of the function indicates that the left end starts at the top of the graph?

OpenStudy (anonymous):

also for the first question I was trying to figure it out but the only opions are f(x − 1) = x3 − 2x2 − x + 3 f(x + 1) = x3 + 4x2 + 3x + 1 f(x) − 1 = x3 + x2 − 2x f(x) + 1 = x3 + x2 − 2x + 2

OpenStudy (abdullah_me):

for the first function, it's f(x+1) that causes the function to shift 1 unit left

OpenStudy (abdullah_me):

im not sure about the second question tho

OpenStudy (anonymous):

ok thank you so much I don't need anymore help :)

OpenStudy (anonymous):

this helped alot thnx